A cone has volume V = 1/3 pi r^2 3
where r is the radius of the base and h is the height. If a cone
has a volume 108pi cm^3 and the radius is 9 centimeters, find its height.
The answer is 4 cm, but could you explain how to find the cone's height?

Answer by rothauserc(4718)

A cone has volume V = 1/3 * pi * r^2 *h
solve equation for h
h = (3*V) / (pi * r^2))
h = (3*108*3.141592654) / (3.141592654 * 81) = 4 cm